# Milan, Italy

### v3

@@ Abstract
Process simulation is gaining attention for its ability to assess potential performance improvements and risks associated with business process changes. The existing literature presents various techniques, generally grounded in process models discovered from event logs or built upon deep learning algorithms. These techniques have specific strengths and limitations. Traditional approaches rooted in process models offer increased interpretability, while those using deep learning excel at generalizing changes across large event logs. However, the practical application of deep learning faces challenges related to managing stochasticity and integrating information for what-if analysis. This paper introduces a novel recurrent neural architecture tailored to discover COnditioned process Simulation MOdels (CoSMo) based on user-based constraints or any other nature of a-priori knowledge. This architecture facilitates the simulation of event logs that adhere to specific constraints by incorporating declarative-based rules into the learning phase as an attempt to fill the gap of incorporating information into deep learning models to perform what-if analysis. Experimental validation illustrates CoSMo’s efficacy in simulating event logs while adhering to predefined declarative conditions, emphasizing both control-flow and data-flow perspectives.
@@ 1 Introduction
Process simulation models are instrumental in a variety of Process Mining (PM) applications including compliance analysis, performance analysis, and what-if analysis. Predominantly, these models are categorized into Data-driven Process Simulation (DDS) models [16, 8, 6, 5], Deep Learning-based (DL) solutions [7, 9], and hybrid approaches combining both [10, 17], each offering unique capabilities and facing distinct challenges. While DDS models excel in simulating control-flow patterns, they are often hindered by oversimplified assumptions about resource behavior and are susceptible to biases from event log characteristics. Conversely, DL models, despite their success in predictive process monitoring [20], struggle with the stochastic nature and the challenges of integrating what-if scenario knowledge, limiting their utility in process simulation [12]. A critical aspect common to both DDS and DL models is the stochastic approach in simulation, relying on randomness to yield realistic outcomes. While this enhances realism, it limits user control over simulated behavior, especially in deep learning approaches, and hinders flexible what-if analysis. Reliance on the most likely sequence often leads to monotonous outcomes, highlighting the need for more control and variability in simulations.
Addressing these challenges and drawing inspiration from related research [14, 11], our contributions comprise (i) a novel conditioned recurrent architecture tailored to capture the relationship between user-defined constraints and the sequential nature of events; (ii) the introduction of a new framework, Conditioned Process Simulation Model (CoSMo), which implements the conditioned recurrent architecture using any type of constraint or a-priori knowledge; (iii) the validation of this approach through experiments considering both control-flow and data-flow perspectives. CoSMo has proven its capability to simulate traces in compliance with constraints represented by the DECLARE language model [19], providing users with significant control over their generative process models and effectively overcoming the stochastic limitations inherent in traditional methods.
@@ 2 Background
A process model $N$ is an abstract representation capturing the sequence and interrelation of activities in historical process executions, as recorded in an event log. It delineates the steps, decision points, and possible paths within a process, aiming to clearly and comprehensively visualize the process flow for analysis, optimization, or comprehension. Process models can be categorized as either imperative or declarative. Imperative models, such as BPMN (Business Process Model and Notation) models [21], define the specific sequential relationships of activities. In contrast, declarative models, like the DECLARE model [19], specify rules and constraints governing the activities without prescribing an exact sequence. On the other hand, process simulation models are defined as a set of techniques that aim at predicting possible behaviors of a process [1]. It might be usually defined as a tuple $M=(N,P)$, where $N$ is a discovered process model (e.g., a BPMN model) and $P$ is a set of parameters needed to define the simulation specifications for the different process perspectives, such as the control-flow, the resource, and time perspectives. These parameters can be manually defined based on expert knowledge [22] or automatically extracted from the event log [8].
@@
Since our contribution makes use of the DECLARE model, we formally describe it here. The process modeling language DECLARE [19] is based on Linear Temporal Logic over finite traces ($\text{LTL}_{f}$[15], which is used to express process semantics. A DECLARE model is defined by a set of activities and constraints on these activities. These constraints are formalized using $\text{LTL}_{f}$ formulae. DECLARE utilizes a subset $\text{LTL}_{f}$ formulae, embodied in predefined patterns known as DECLARE templates. Thus, each template is an $\text{LTL}_{f}$ formula containing placeholders, which are replaced by specific activities to form a constraint. For example, considering the template $ExclusiveChoice(a,b)$, where $a$ and $b$ are arbitrary activities, it is stipulated that either $a$ or $b$ must occur at least once, but not both. See [11, 3] for more details on the $\text{LTL}_{f}$ semantics. More specifically, given a trace $t$ of length $s$ and an $\text{LTL}_{f}$ formula $\phi$, the satisfaction of $\phi$ in $t$ at position $i$ is denoted by $t,i\models\phi$. The notation $t\models\phi$ is used as a shortcut to indicate that the formulae $\phi$ holds over the entire trace $t$ starting from the beginning. A set of constraints $C=\{\phi_{0},...,\phi_{m}\}$ is a model for a log $L$, denoted by $L\models C$, and $\phi$ is satisfiable if it admits at least one trace $t$ such that $t\models\phi_{0}\land...\land\phi_{m}$ for each $t\in L$. Thus, a DECLARE model can be formalized as the set of $\text{LTL}_{f}$ formulae $C$ expressing its constraints. Among all the existing $\text{LTL}_{f}$ formulae, DECLARE selects predefined templates, which can be gathered into four main groups according to their semantics. Given the arbitrary activities $a$ and $b$, the templates can be defined as follows: Existence ($\mathcal{E}$): given $a$, this group of templates checks the number of occurrences, its absence, or its position in the trace; Choice ($\mathcal{C}$): given $a$ and $b$, this group of templates check if $a$ or $b$ occurs, or if $a$ and $b$ occur together; Positive Relations($\mathcal{PR}$): given $a$ and $b$, this group checks the relative position of the activities. For instance, if $a$ occurs, $b$ eventually occurs; Negative Relations($\mathcal{NR}$): given $a$ and $b$, this group checks if the activities do not occur together or do not occur in a certain order. For instance, if $a$ occurs, $b$ never occurs. Following the notations presented by Donadello et al. [11], the complete set of DECLARE templates is represented as $\mathcal{DT}=\mathcal{E}\cup\mathcal{C}\cup\mathcal{PR}\cup\mathcal{NR}$. Due to lack of space, we are not describing all the templates of $\text{LTL}_{f}$ formulae, but all semantics outlined by the authors (see Table 2 in  [11]) were utilized in this work.
@@ 3 Current State and Limitations of Process Simulation
DDS models are built upon process models discovered from event logs [22, 4, 5]. This group of simulation models is the most popular in the current literature and includes several proposals such as the PGL2 [4] and SIMOD [8]. These methods are typically designed to simulate different control-flow patterns by manipulating user-based requirements such as the number of gates and the amount of noise. However, DDS models are constrained by oversimplified assumptions about resource behavior, neglecting complexities such as multitasking, batching, and resource sharing [13]. In addition, process discovery algorithms are biased by event log and discovery algorithm characteristics, meaning that they may behave differently depending on the scenario [25].
Due to the outstanding performance of deep learning models in predictive process monitoring [24, 14, 18, 7, 20], recent efforts have been devoted to exploring their potential in process simulation [9, 10, 17]. However, DL models are problematic in the context of process simulation due to their stochastic nature and the difficulty of incorporating what-if scenario information into the DL model [9]. These limitations are quite relevant for process simulation since the user is not able to control the simulation and the model is not able to learn how to simulate possible process changes imposed by the user. For this reason, deep learning models have been limited to working as generative models [9] or have been combined with PM techniques to overcome these limitations [10, 17].
A common characteristic among process simulation models is the stochasticity inserted during simulation. In general, the simulation process requires a sampling step to draw information from the probability distributions, and applying a certain level of randomness helps models produce more realistic simulations [1]. For example, DDS-based simulations are supported by statistical information extracted from the event logs, such as branch probabilities and activity duration time distributions. On the other hand, DL-based approaches consist of learning the underlying data distributions and drawing event attributes from the probability distributions returned by the learned model. Although this stochastic approach makes simulations more realistic [1, 7], the complete dependence on randomness limits the flexibility of the user to control the simulated behaviors. Furthermore, despite the efforts to incorporate a-priori knowledge or inter-case features to improve the predictive performance [14, 11], the incorporation of information for what-if analysis is still challenging for DL-based solutions [12].
To illustrate this stochastic problem, let us consider the problem of next activity prediction in PM [20]. The goal here is to always determine the most likely next activity in an ongoing process. Relying solely on the most likely sequence results in monotonous, identical trace variants. Typically, this problem is addressed by sampling the next activity from a probability distribution, using techniques such as multinomial sampling [24] or beam search [14]. However, such stochastic methods reduce the user’s ability to precisely control the simulation, especially when performing what-if analyses.
@@ 4.1 Conditioned Recurrent Architecture
Our proposed architecture starts by taking a sequence of events $x$ as input, where each event element is composed of different attributes, as described in Section 2. Alongside this sequence, the architecture is introduced with external feature vectors $c$ representing a set of constraints. We extend the vanilla recurrent neural network (RNN) by introducing a new learnable parameter $Q$ to capture the relation between the condition $\Phi$ and the sequential input, followed by a residual connection between each recurrent cell:
$\begin{gathered}h_{j}=f(Ux_{j}+Wh_{j-1}+Q\Phi)\\ y_{j}=g(Vh_{j})+x_{j}\end{gathered}$ (1)
Where $h_{j}$ is the hidden state at time step $j$. This represents the “memory” of the network up to that point in time. On the other hand, $U$, $W$, and $V$ are the learnable parameters from the vanilla RNN, and each plays a distinct role in transforming the input and hidden states. The parameter $Q$ specifically learns how the constraint feature vector $\Phi$ relates to the input sequence at every timestep. The key for the proposed recurrent network is that these condition vectors are external to the primary sequence but highly relevant to the task at hand. Finally, $g$ is a non-linear activation function, essential for enabling the network to capture complex patterns and the residual connection at $y_{j}$ helps to improve and stabilize the learning phase. This residual technique is common in other deep learning applications and modern recurrent architectures, such as LSTMs, have a similar characteristic (e.g., the forgetting gate mechanisms). The bias terms are omitted for simplicity. We depict the new architecture in Figure 0(b).
@@ 5.1 Experimental Setup
Simulation. Building upon the methodology outlined in Section 4, we initiate by selecting the constraints $\Phi_{\text{test}}$, represented in our study as groups of DECLARE templates, each comprising a set of $\text{LTL}_{f}$ rules. To condition the simulation, we then identify a subset $\Phi^{\prime}$ from these constraints. Ensuring consistency and fairness in our evaluation across all event logs, we uniformly select the same DECLARE templates. As previously mentioned, these rules are expressed as vectors of binary values, where ‘1’ indicates rule fulfillment, and ‘0’ denotes non-fulfillment. Our primary aim is to invert one specific subset of rules to ‘enforce’ desired process behavior. However, since we are using all the available DECLARE templates, we must ensure that the rules are not contradictory. For instance, inverting the ‘existence’ rule while maintaining the ‘absence’ rule could create contradictory conditions, leading to a paradox in the model’s inference process. We perform a careful selection of templates to avoid any inconsistencies: we do not include the $\mathcal{NR}$ template group in our evaluation, as it is inversely related to the $\mathcal{PR}$ group, in general. Such meticulous template selection ensures coherence and efficacy in our model’s simulation capabilities. Thus, if we want to ensure the ‘existence’ rule, the ‘absence’ must be set as the opposite as well. Therefore, for each group of templates, we pick the following rules as references: the Existence(a) regarding the group $\mathcal{E}$, the Exclusive choice(a,b) regarding the group $\mathcal{C}$, and the Chain Response(a, b) regarding the group $\mathcal{PR}$. Subsequently, the following rules are also picked to avoid the aforementioned inconsistency issue: (i) for group $\mathcal{E}$, the templates Absence(a) and Exactly(1, a) are chosen; (ii) for group $\mathcal{C}$, no extra template is selected; and (iii) for group $\mathcal{PR}$, we select the Alternate Response(a, b), Precedence(a, b), and Response(a, b) templates. The remaining rules are preserved ($\Phi^{p}$) and the picked ones are inverted ($\Phi^{\prime}$) to compose the final set of constraints $\Phi^{s}=\Phi^{p}\cup\Phi^{\prime}$ to be simulated. Each of these templates and its details is described in [11]. We selected the activities based on their frequencies (we picked the 3rd quartile), and the activities incorporated into each group of templates for each event log are described in Table 2.
@@
Evaluation and Validation. Upon generating a simulated event log under these inverted constraints, we employ conformance checking to verify the fulfillment of the modified rule set. In this evaluation, we split the traces between two classes to quantify the effectiveness of the simulation: one where the reference rule must be fulfilled and the other where it must not. For instance, a trace where the reference rule must be fulfilled (e.g., Existence(a)=1) will be a true positive if the simulation conforms to the rule, otherwise, it will be a false negative. The performance is then reported through the calculation of weighted precision, a metric chosen specifically for its ability to provide nuanced insight into the degree to which the simulated traces adhere to or deviate from their intended behavior. The use of weighted precision is particularly appropriate in this context because it allows for a nuanced assessment of each trace’s adherence to the imposed rules, recognizing the varying importance of different rules within the process, thereby providing a more granular and meaningful measure of CoSMo’s effectiveness and ability to replicate or alter specific process behaviors as needed. In addition, we include in this evaluation two baselines against the aforementioned inverted subset. First, we compare CoSMo with this perturbed set of constraints ($\Phi^{s}$) as input to CoSMo with the original set of constraints ($\Phi$) as input, which is the simplest way to assess whether the model actually changes or maintains the input set of rules. Simulating the Original and Inverted subset of rules is an alternative to performing an ablation study, which is a common practice in the machine learning literature to assess the impact of a particular component in a model. Second, we compare CoSMo to a Vanilla simulation without any constraints, which is the most common way to stochastically generate processes using deep learning in the literature [9]. The intuition behind this comparison is to assess whether the model is actually learning the constraints and whether the constraints are affecting the simulation. We use the architecture introduced in [24] as a vanilla baseline, since it overcomes other approaches from the predictive monitoring literature [20]. We run each simulation 10 times to ensure the robustness of the results since the stochastic nature of the model can lead to different results in each run.

### v4

@@ Abstract
Process simulation is gaining attention for its ability to assess potential performance improvements and risks associated with business process changes. The existing literature presents various techniques, generally grounded in process models discovered from event log data or built upon deep learning algorithms. These techniques have specific strengths and limitations. Traditional data-driven approaches offer increased interpretability, while deep learning-based excel at generalizing changes across large event logs. However, the practical application of deep learning faces challenges related to managing stochasticity and integrating information for what-if analysis. This paper introduces a novel recurrent neural architecture tailored to discover COnditioned process Simulation MOdels (CoSMo) based on user-based constraints or any other nature of a-priori knowledge. This architecture facilitates the simulation of event logs that adhere to specific constraints by incorporating declarative-based rules into the learning phase as an attempt to fill the gap of incorporating information into deep learning models to perform what-if analysis. Experimental validation illustrates CoSMo’s efficacy in simulating event logs while adhering to predefined declarative conditions, emphasizing both control-flow and data-flow perspectives.
@@ 1 Introduction
Process simulation models are instrumental in a variety of Process Mining (PM) applications including compliance analysis, performance analysis, and what-if analysis. Predominantly, these models are categorized into Data-driven Process Simulation (DDS) models [16, 8, 6, 5], Deep Learning-based (DL) solutions [7, 9], and hybrid approaches combining both [10, 17], each offering unique capabilities and facing distinct challenges. While DDS models excel in simulating control-flow patterns, they are often hindered by oversimplified assumptions about resource behavior and are susceptible to biases from event log characteristics. Conversely, DL models, despite their success in predictive process monitoring [20], struggle with the stochastic nature and the challenges of integrating what-if scenario knowledge, limiting their utility in process simulation [12]. A natural aspect common to both DDS and DL models is the stochastic approach in simulation, relying on randomness to yield realistic outcomes [1]. While this enhances realism, it might limit user control over simulated behavior and hinder practical flexibility (e.g., for what-if analysis). In this sense, DDS models overcome this issue by providing tunable parameters [8]. On the other hand, DL-based process simulation models typically do not provide this flexibility. These models are often seen as black boxes, with no options for users to influence the simulation results, relying solely on random sampling methods to draw predictions from the probability distributions outputted by the model.
Therefore, in this paper, we motivate the use of deep learning for process simulation due to its successful accomplishments in other process mining tasks, and we address the problem associated with the lack of flexibility of current DL-based solutions for process simulations [12]. By proposing a novel methodology to introduce user-directed controls and guidance within these models, our work advances the fields by introducing a new bridge between advanced process simulation capabilities using deep learning and practical usability needs. Thus, we comprise our contribution as follows: (i) a novel conditioned recurrent architecture tailored to capture the relationship between user-defined constraints and the sequential nature of events; (ii) the introduction of a new framework, Conditioned Process Simulation Model (CoSMo), which implements the conditioned recurrent architecture using any type of constraint or a-priori knowledge; (iii) the validation of this approach through experiments considering both control-flow and data-flow perspectives. CoSMo has proven its capability to simulate traces in compliance with constraints represented by the DECLARE language model [19], providing users with significant control over their generative process models and effectively overcoming the stochastic limitations inherent in traditional DL-based process simulation models.
@@ 2 Background
A process model $N$ is an abstract representation capturing the sequence and interrelation of activities in historical process executions, as recorded in an event log. Process models can be classified as either imperative, like BPMN [21], which define specific relationships of activities, or declarative, like the DECLARE model [19], which set rules and constraints without specifying exact sequences. On the other hand, process simulation models are defined as a set of techniques that aim at stochastically predicting possible behaviors of a process [1]. It might be usually defined as a tuple $M=(N,P)$, where $N$ is a discovered process model (e.g., a BPMN model) and $P$ is a set of parameters needed to define the simulation specifications for the different process perspectives, such as the control-flow, the resource, and time perspectives. These parameters can be manually defined based on expert knowledge [22] or automatically extracted from the event log [8].
@@
We now provide essential definitions related to the DECLARE model used in this work. The process modeling language DECLARE [19] is based on Linear Temporal Logic over finite traces ($\text{LTL}_{f}$[15], which is used to express process semantics. A DECLARE model is defined by a set of activities and constraints on these activities. These constraints are formalized using $\text{LTL}_{f}$ formulae. DECLARE utilizes a subset $\text{LTL}_{f}$ formulae, embodied in predefined patterns known as DECLARE templates. Following the notations presented by Donadello et al. [11], we define the complete set of DECLARE templates as $\mathcal{DT}=\mathcal{E}\cup\mathcal{C}\cup\mathcal{PR}\cup\mathcal{NR}$, where each template is defined as follows. Existence ($\mathcal{E}$): given an arbitrary activity $a$, this group of templates checks the number of occurrences, its absence, or its position in the trace; Choice ($\mathcal{C}$): given $a$ and $b$, this group of templates check if $a$ or $b$ occurs, or if $a$ and $b$ occur together; Positive Relations($\mathcal{PR}$): given $a$ and $b$, this group checks the relative position of the activities. For instance, if $a$ occurs, $b$ eventually occurs; Negative Relations($\mathcal{NR}$): given $a$ and $b$, this group checks if the activities do not occur together or do not occur in a certain order. For instance, if $a$ occurs, $b$ never occurs. In the context of this work, we omit further details and formal definitions as we follow the same as defined in the literature [11, 3].
@@ 3 Current State and Limitations of Process Simulation
DDS models are built upon process models discovered from event logs [22, 4, 5], and popular proposals include PGL2 [4] and SIMOD [8]. These methods are typically designed to simulate different control-flow patterns by manipulating user-based requirements such as the number of gates and the amount of noise. However, DDS models are constrained by oversimplified assumptions about resource behavior, neglecting complexities such as multitasking, batching, and resource sharing [13]. Despite the outstanding performance of deep learning models in predictive monitoring [24, 14, 18, 7, 20], its use in process simulation is limited by its stochastic nature and difficulty in adapting to what-if scenarios [9], often leading to the limited of DL as generative models [9] or hybrid models with process mining techniques [10, 17].
A common characteristic among process simulation models is the stochasticity inserted during simulation. The simulation process requires a sampling step to draw information from the probability distributions, and applying a certain level of randomness helps models produce more realistic simulations [1]. For example, DDS-based simulations are supported by statistical measures extracted from the event logs, such as branch probabilities and activity duration time distributions. On the other hand, DL-based approaches consist of learning the underlying data distributions and drawing event attributes from the probability distributions returned by the learned model. Although this stochastic approach makes simulations more realistic [1, 7], the complete dependence on randomness limits the flexibility of the user to control the simulated behaviors [12].
To illustrate this stochastic problem, let us consider the problem of next activity prediction in PM [20]. The goal here is to always determine the most likely next activity in an ongoing process. Conversely, in process generation or simulation, diversity is key because realistic process traces typically vary. Relying solely on the most likely sequence results in monotonous, identical trace variants. Typically, this problem is addressed by sampling the next activity from a probability distribution, using techniques such as multinomial sampling [24] or beam search [14]. However, such stochastic methods reduce the user’s ability to precisely control the simulation, especially when performing what-if analyses.
@@ 4.1 Conditioned Recurrent Architecture
Our proposed architecture starts by taking a sequence of events $x$ as input, where each event element is composed of different attributes, as described in Section 2. Alongside this sequence, the architecture is introduced with external feature vectors $c$ representing a set of constraints. We extend the vanilla recurrent neural network (RNN) by introducing a new learnable parameter $K$ to capture the relation between the condition $\Phi$ and the sequential input, followed by a residual connection between each recurrent cell:
$\begin{gathered}h_{j}=f(Ux_{j}+Wh_{j-1}+K\Phi)\\ y_{j}=g(Vh_{j})+x_{j}\end{gathered}$ (1)
Where $h_{j}$ is the hidden state at time step $j$. This represents the “memory” of the network up to that point in time. On the other hand, $U$, $W$, and $V$ are the learnable parameters from the vanilla RNN, and each plays a distinct role in transforming the input and hidden states. The parameter $K$ specifically learns how the constraint feature vector $\Phi$ relates to the input sequence at every timestep. The key for the proposed recurrent network is that these condition vectors are external to the primary sequence but highly relevant to the task at hand. In simple terms, a basic RNN learns a hidden state for each event in the sequence. Our model introduces extra parameters, $K$, which act like an additional hidden state. This extra state learns the relationship between the set of rules at each time step and its corresponding hidden state. Finally, $g$ is a non-linear activation function, essential for enabling the network to capture complex patterns and the residual connection at $y_{j}$ helps to improve and stabilize the learning phase. This residual technique is common in other deep learning applications and modern recurrent architectures, such as LSTMs, have a similar characteristic (e.g., the forgetting gate mechanisms). The bias terms are omitted for simplicity. We depict the new architecture in Figure 1(b).
@@ 5.1 Experimental Setup
Simulation. Building upon the methodology outlined in Section 4, we initiate by selecting the constraints $\Phi_{\text{test}}$, represented in our study as groups of DECLARE templates. To condition the simulation, we then identify a subset $\Phi^{\prime}$ from these constraints to be inverted. Given the frequency distribution of the set of rules, a pair of activities (or a single activity when the template is the Existence) is drawn from it taking into consideration only the rules lying between the 1st and 3rd quantile of that distribution. By setting this boundary, we aim at not selecting rules with low variance in order to accomplish greater generalization. Despite the arbitrary intuition behind this selection, in the sense of business-driven choices, we propose this data-driven approach to ensure a fair evaluation across all event logs and eliminate the need for an expert to define the rules to be simulated for each event log. Although this is a clear limitation in real-world scenarios, it effectively validates our scientific methodology.
Subsequently, in order to ensure consistency in our evaluation across all event logs, we uniformly select the same DECLARE templates. As previously mentioned, these rules are expressed as vectors of binary values, where ‘1’ indicates rule fulfillment, and ‘0’ denotes non-fulfillment. Our primary aim is to invert one specific subset of rules to ‘enforce’ desired process behavior. However, since we are using all the available DECLARE templates, we must ensure that the rules are not contradictory. For instance, inverting the ‘existence’ rule while maintaining the ‘absence’ rule could create contradictory conditions, leading to a conflict in the model’s inference process. We perform a careful selection of templates to avoid any inconsistencies: we do not include the $\mathcal{NR}$ template group in our evaluation, as it is inversely related to the $\mathcal{PR}$ group, in general. Such meticulous template selection ensures coherence and efficacy in our model’s simulation capabilities. Thus, if we want to ensure the ‘existence’ rule, the ‘absence’ must be set as the opposite as well.
Therefore, for each group of templates, we pick the following rules as references: the Existence(a) regarding the group $\mathcal{E}$, the Exclusive choice(a,b) regarding the group $\mathcal{C}$, and the Chain Response(a, b) regarding the group $\mathcal{PR}$. Subsequently, the following rules are also picked to avoid the aforementioned inconsistency issue: (i) for group $\mathcal{E}$, the templates Absence(a) and Exactly(1, a) are chosen; (ii) for group $\mathcal{C}$, no extra template is selected; and (iii) for group $\mathcal{PR}$, we select the Alternate Response(a, b), Precedence(a, b), and Response(a, b) templates. The remaining rules are preserved ($\Phi^{p}$) and the picked ones are inverted ($\Phi^{\prime}$) to compose the final set of constraints $\Phi^{s}=\Phi^{p}\cup\Phi^{\prime}$ to be simulated. Each of these templates and its details is described in [11]. We selected the activities based on their frequencies (we picked the 3rd quartile), and the activities incorporated into each group of templates for each event log are described in Table 2.
@@
Evaluation and Validation. Upon generating a simulated event log under these inverted constraints, we employ conformance checking to verify the fulfillment of the modified rule set. In this evaluation, we split the traces between two classes to quantify the effectiveness of the simulation: one where the reference rule must be fulfilled and the other where it must not. For instance, a trace where the reference rule must be fulfilled (e.g., Existence(a)=1) will be a true positive if the simulation conforms to the rule, otherwise, it will be a false negative. The performance is then reported through the calculation of weighted precision, a metric chosen specifically for its ability to provide nuanced insight into the degree to which the simulated traces adhere to or deviate from their intended behavior. The use of weighted precision is particularly appropriate in this context because it allows for a nuanced assessment of each trace’s adherence to the imposed rules, recognizing the varying importance of different rules within the process, thereby providing a more granular and meaningful measure of CoSMo’s effectiveness and ability to replicate or alter specific process behaviors as needed.
In addition, we include in this evaluation two baselines against the aforementioned inverted subset. First, we compare CoSMo with this perturbed set of constraints ($\Phi^{s}$) as input to CoSMo with the original set of constraints ($\Phi$) as input, which is the simplest way to assess whether the model actually changes or maintains the input set of rules. Simulating the Original and Inverted subset of rules is an alternative to performing an ablation study, which is a common practice in the machine learning literature to assess the impact of a particular component in a model. In practical terms, by providing the Original set of rules, we expect to validate the model’s capability to learn the relationship between those rules and the process semantics. On the other hand, the Inverted subset allows us to validate the model’s capabilities to generalize. More specifically, this approach could be thought of as the changes that we aim to do in a process, which could characterize a what-if scenario. Second, we compare CoSMo to a Vanilla simulation without any constraints, which is the most common way to stochastically generate processes using deep learning in the literature [9]. The intuition behind this comparison is to assess whether the model is actually learning the constraints and whether the constraints are affecting the simulation. We use the architecture introduced in [24] as a vanilla baseline, since it overcomes other approaches from the predictive monitoring literature [20]. We run each simulation 10 times to ensure the robustness of the results since the stochastic nature of the model can lead to different results in each run.